OpenStudy (anonymous):

A store owner purchases 30 boxes of flash drives. Of those boxes 1/5 are given to the employees, and the rest are placed on the shelves to be sold. Each employee receives 2/3 of the boxes that are given to the staff. How many employees are there?

OpenStudy (anonymous):

A. 4 employees B. 9 employees C. 12 employees D. 15 employees

OpenStudy (anonymous):

Assuming "staff" and "employees" are synonymous, this question makes absolutely no sense. Typo?? How can everybody receive 2/3 of the boxes given to themselves???

OpenStudy (anonymous):

Hey lets get a pizza! Everybody gets 2/3 of the pizza! Huh?!?

OpenStudy (texaschic101):

1/5 of 30 = 30/5 = 6 boxes of flashdrives are given to the employees. 2/3 of 6 = 12/3 = 4 There are 6 boxes and each employee receives 2/3 of each box. Making 4 employees. You might get a second opinion on this....it is kinda confusing.

OpenStudy (anonymous):

Ikr

OpenStudy (anonymous):

It is part of my sister's math lesson today, my mom and I both think it is worded weirdly. The thing is if 2/3 of the 6 boxes go to each employee, then that is 4 boxes per employee and there would be two boxes left over. That would mean there are 1 and 1/2 employees. That just makes no sense.

OpenStudy (anonymous):

ahh..! i think they mean each employee receives 2/3 of 1 box!!! 6 boxes divided by 2/3 = 9 There are 9 employees.

OpenStudy (texaschic101):

that is probably right

OpenStudy (anonymous):

it shouldn't read "2/3 of the boxes" it should read "2/3 of a box"

OpenStudy (texaschic101):

Badly worded problem

OpenStudy (anonymous):

Yeah the whole curriculum is like that

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

thank Bangkok....he figured it out :)

OpenStudy (anonymous):

You still tried to help though, and I appreciate it.